OpenStudy (anonymous):

x = -9 a. x-intercept is ( -9, 0) No y-intercept c. No x-intercept y-intercept is (0, -9) b. No x-intercept y-intercept is ( -9, 0) d. x-intercept is (0, -9) No y-intercept

OpenStudy (idku):

What do you think ?

OpenStudy (idku):

You can eliminate some of the choices without even looking at the question.

OpenStudy (anonymous):

its either a or d. im thinking its D i just want to make sure

OpenStudy (idku):

D ? Have you read choice D ?

OpenStudy (idku):

x intercept is written as (a,0) and not as (0,a) y intercept is written as (0,a) and not as (a,0)

OpenStudy (anonymous):

oh so it would be A.

OpenStudy (idku):

Yes, exactly !

OpenStudy (anonymous):

y = x + 8 a. x-intercept is (0, 8) y-intercept is ( -8, 0) c. x-intercept is ( -8, 0) y-intercept is (0, 8) b. x-intercept is (0, -8) y-intercept is ( 8, 0) d. x-intercept is ( 8, 0) y-intercept is (0, -8)

OpenStudy (anonymous):

what about this one

OpenStudy (idku):

You have a line y = x + 8 To find the y-intercept set x=0 and solve for y. You do it like this, y = x + 8 y = (0) + 8 y= ? You tell me remember that y intercept v would be written as (0,v) To find the x-intercept set y=0 and solve for y. You do it like this, y = x + 8 (0) = x + 8 x= ? You tell me

OpenStudy (idku):

remember that x intercept v would be written as (v,0)

OpenStudy (anonymous):

the answer is d?

OpenStudy (anonymous):

its either d or c

OpenStudy (anonymous):

and if you could help me through this it would mean a lot to me, i dont understand it at all y = 7x + 43 y = 2x + 13 a. ( -6, 1) c. ( 6, 1) b. ( 6, -1) d. No solution

OpenStudy (anonymous):

@idku

OpenStudy (anonymous):

@AkashdeepDeb @anonymous_user @alakara_kibo @alyssajobug @AmorVida @ashhh4356787 @brei @Bsho1997 @BrianMatthew @bbblakeee @ballinbro @cookiibabii93 @chelseamonae @DUDE..IM..A..DUCK @deezgurls @Drsuz98 @DUDE..LOOK..ITS..SGTC @dg2 @desperatemom @dg2 @geekfromthefutur @hazyskies @fireheartroo

OpenStudy (anonymous):

if i cant get help , nobody can

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

well first what do you think?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

dont you have to subtract

OpenStudy (anonymous):

kinda. what you need to do is to change one of the equations so that it is x=y....

OpenStudy (anonymous):

i just need the answer really quick

OpenStudy (anonymous):

(-6,1) you just change on them to that so that you can plug it into the other